PAir=1.29 kg/m^3 and PHelium= 0.179 kg/m^3

Then what volume of helium is needed if a baloon is to lift a load of 180 kg (incliding the weight of the empty balloon)?
Answer #1

Let V is the require volume of helium

mass of helium = density*volume

= 0.179*V

We know, Buoynat force = weight of the dislpced fluid

here, Buyont force = weight 180kg mass + weight of helium

so,weight of the dislpced fluid = weight 180kg mass + weight of helium


rho_air*V*g = 180*g + rho_helium*V*g

rho_air*V = 180 + rho_helium*V


V*(rho_air - rho_helium) = 180

V = 180/(rho_air - rho_helium)

= 180/(1.29 - 0.179)

= 162 m^3 <<<<<<<------------------Answer